OpenStudy (anonymous):

find the value of b such that the system of equations bx-15y=-20 4x+10y=-17 does not have a solution. what is b=???

OpenStudy (anonymous):

what is the condition for two lines not to have a solution?

OpenStudy (anonymous):

b = - 6 ?

OpenStudy (anonymous):

not sure what your question is. but what is b? that's all i need to know. and how to find it. I've done this long time ago can't remember how to do these.

OpenStudy (anonymous):

nice going dhatra :)i like ur approach :)

OpenStudy (anonymous):

well, two lines are said to have a solution at the point of their intersection. if the two lines are parallel, then they will not intersect, and hence will not have a solution. two lines are parallel if 1. they have the same slope 2. they are separated by some distance.

OpenStudy (anonymous):

why thank you, brackett. :)

OpenStudy (anonymous):

dhatraditya. Explain to me, how would you solve for the variable b?

OpenStudy (anonymous):

first lets rearrange each equation to the form y = mx+b

OpenStudy (anonymous):

@beehive: for the equations to hav no solution, the lines must be parallel and hence their slope has to be same...therefore u hav b/15=-4/10 and so u get, b=-6

OpenStudy (anonymous):

yup, as brackett said.

OpenStudy (anonymous):

brackett how did you get here? b/15=-4/10 ????

OpenStudy (anonymous):

i equated both the slopes

OpenStudy (anonymous):

@ beehive: rearrange both the equations in the slope intercept form, ie y = mx+b and post what you get.

OpenStudy (anonymous):

do you know how to rearrange the equations, beehive?

OpenStudy (anonymous):

yes i do. hold on a sec. I'm trying to get this straight in my head. i've done this like 4 years ago. forgot it all.

OpenStudy (anonymous):

alright, whenever you are ready

OpenStudy (anonymous):

for the first one. i got \[y=\frac{-bx}{-15}+\frac{4}{3}\] the second one i got. \[y=-\frac{2}{5}x-\frac{17}{10}\] this seem right?

OpenStudy (anonymous):

right. that is correct. now, for two equations expressed in the slope intercept form to be parallel, their slopes must be equal. so -b/-15 = -2/5

OpenStudy (anonymous):

well i understand you can rewrite the y=(-bx)/(15) to \[y=\frac{-b}{15}x\]

OpenStudy (anonymous):

ok i see. Thanks.

OpenStudy (anonymous):

no, -bx/-15 = -1*bx/-1*15 = bx/15

OpenStudy (anonymous):

you are welcome